Calculate the magnitudes of manetic intensity H and the magnetic field B at the centre of a
15oo - turn solenoid which is 0.22m long and carries a current of 1.5A (UO= 4×3.14×10¯7Hm¯¹
1
Expert's answer
2020-04-13T10:21:34-0400

The magnetic field at the center of solenoid

"B=\mu_0\frac{NI}{L}""=4\pi\times 10^{-7}\times \frac{1500\times 1.5}{0.22}=0.013\:\rm T"

The  magnitudes of magnetic intensity at the center of solenoid

"H=B/\mu_0=0.013/(4\pi\times 10^{-7})=10227\:\rm A/m."